OpenStudy (anonymous):

You have $100 saved and a plan to save $20 each month. Write a rule for the amount saved (in dollars) as a function of the number of months from now. x | y 0 | 100 1 | 120 2 | 140 3 | 160 4 | 180 5 | 200 6 | 220 7 | 240 8 | 260 9 | 280 10| 300 11| 320 12| 340

OpenStudy (anonymous):

y = 20x + 100

OpenStudy (anonymous):

Because you already have 100 dollars that will be your starting point (or your y-intercept). Then you are saving 20 dollars every month. "x" stands for the number of months. So for example: Say you have saved up for 4 months so far. You would plug 4 into the x value: y = 20x + 100 y = 20(4) + 100 y=80 +100 y= 80 (you can check your chart for confirmation.)